# Edge Importance in Complex Networks

### v1

@@ Abstract
Keywords: Network analysis, Sensitivity analysis, Edge importance
@@ 1 Introduction
Networks are helpful for modeling complex interactions between entities. A network can be represented by a graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E},\mathcal{W}\rangle$, which consists of a set of vertices or nodes $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, a set of edges =$\mathcal{E}=\{e_{1},e_{2},\dots,e_{m}\}$ that connect the vertices, and a set of nonnegative edge weights $\mathcal{W}=\{w_{11},w_{21},\ldots,w_{nn}\}$. The weight $w_{ij}$ is positive if there is an edge pointing from vertex $v_{i}$ to vertex $v_{j}$; while $w_{ij}=0$ signifies that there is no such edge. Edges may be directed (and then model one-way streets) or undirected (and then model two-way streets). If the graph models a road network in which the vertices model intersections and the edges model roads, then the weights may, e.g., be proportional to the amount of traffic along each road. A network in which all positive weights are one is said to be unweighted. Descriptions and many applications of networks are provided by, e.g., Estrada [11] and Newman [19].
@@
Let $\mathbf{A}^{k}=[a_{ij}^{(k)}]_{i,j=1}^{n}$ for $k=1,2,\ldots~{}$, where $a_{ij}^{(1)}=w_{ij}$ for $1\leq i,j\leq n$. A nonvanishing entry $a_{ij}^{(k)}$ for some $k>0$ indicates that there is at least one walk of $k$ edges from vertex $v_{i}$ to vertex $v_{j}$. The denominators in the terms of the expansion (1) ensure that the expansion converges and that terms $\frac{\mathbf{A}^{k}}{k!}$ with $k$ large contribute only little to $\exp(\mathbf{A})$. It follows that short walks typically are more important than long ones, which is in agreement with the intuition that messages propagate better along short walks than along long ones. This led Estrada and Rodriguez-Velazquez [13] to use $\exp(\mathbf{A})$ to study properties of a graph; we use the function $\exp_{0}(\mathbf{A})$ because the term with the identity matrix $\mathbf{I}$ in the expansion of $\exp(\mathbf{A})$ has no natural interpretation in the context of network modeling. Other functions also can be used such as a resolvent or a Mittag-Leffler function; see Estrada and Higham [12] and Arrigo and Durastante [2] for discussions.
Estrada and Rodriguez-Velazquez [13] define for graphs with a symmetric adjacency matrix the communicability matrix $\exp(\mathbf{A})$; we will use the matrix
@@
The entry $c_{ij}$, $i\neq j$, is referred to as the communicability between the vertices $v_{i}$ and $v_{j}$. A relatively large value implies that it is easy for the vertices $v_{i}$ and $v_{j}$ to communicate. Estrada and Rodriguez-Velazquez [13] measure the importance of the vertex $v_{i}$ of an undirected graph by the subgraph centrality $c_{ii}+1$; we will use $c_{ii}$. Related measures of communicability can be defined when the adjacency matrix $\mathbf{A}$ is nonsymmetric; see [6].
@@
where $e=[1,1,\ldots,1]^{T}\in\mathbb{R}^{n}$ denotes the vector with all entries $1$ and the superscript T stands for transposition. Benzi and Klymko [3] introduced the related measure $e^{T}\exp(\mathbf{A})e$, which differs from (2) by the additive constant $n$. Note that the expression (2) is invariant under transposition. We have
@@
We are interested in investigating the importance of the edges of a graph ${\mathcal{G}}$ and, in particular, in determining which edge weights can be reduced or set to zero without significantly affecting the total communicability. A possible approach to investigate the importance of edges is to consider the line graph associated with the graph. The edges of ${\mathcal{G}}$ correspond to vertices in the associated line graph, and the importance of the edges in ${\mathcal{G}}$ can be measured by the subgraph centrality of the vertices of the line graph. This approach is investigated in [7, 8]. However, it is quite cumbersome to construct the line graph except for small graphs. A simple heuristic technique was proposed by Arrigo and Benzi [1], who define the importance of an edge in terms of the importance of the vertices where the edge starts and ends. However, this approach is not guaranteed to correctly rank the importance of edges; see [7, Section 5.4] for an example. More recently an approach that uses the right and left Perron vectors of the adjacency matrix in combination with the Wilkinson perturbation to determine which weights to increase in order to increase the total communicability has been described in [4, 21]. An analogous technique is applied in [10] to discern which weights in a weighted multilayer network can be decreased without affecting the total communicability significantly.
Another approach to study the importance of edges is to evaluate the Fréchet derivatives of the total communicability (2) with respect to the weights. This approach was advocated by De la Cruz Cabrera et al. [4] and recently Schweitzer [27] described how to speed up the computations. Introduce the gradient
@@
This paper is organized as follows. Section 2 discusses how the gradient can be applied to assess the importance of edges. The first part of the section is concerned with small to medium-sized problems for which it is feasible to evaluate the gradient (3). The latter part of the section discusses the application of Krylov subspace methods to project large-scale problems to problems of fairly small size. The computations use a result by Schweitzer [27] on the evaluation of Fréchet derivatives, but differ in various aspects. Section 3 reviews methods described in [4, 10, 21] based on evaluating the right and left Perron vectors and the Wilkinson perturbation to determine important and unimportant edges. Numerical examples are reported in Section 4. These examples compare the methods of Sections 2 and 3. Section 5 contains concluding remarks.
We conclude this section with comments on some related methods. A scheme that combines regression, soft-thresholding, and projection is applied in [5] to approximate an unweighted network by a simpler unweighted network. This scheme performs well but may be expensive and is restricted to unweighted networks. Massei and Tudisco [17] consider the problem of determining a low-rank perturbation $\mathbf{E}\in\mathbb{R}^{n\times n}$ to the adjacency matrix $\mathbf{A}$ so that the perturbed matrix $\mathbf{A}+\mathbf{E}$ maximizes the robustness of the network. For instance, $\mathbf{E}$ may be chosen to maximize the trace of $f(\mathbf{A}+\mathbf{E})-f(\mathbf{A})$ for a user-specified matrix-valued function $f$. The perturbation $\mathbf{E}$ is determined by a greedy algorithm for solving an optimization problem. A careful comparison with this method is outside the scope of the present paper.
@@ 2.1 Methods for small to medium-sized networks
Let the matrix function $f:\mathbf{A}\in\mathbb{R}^{n\times n}\to f(\mathbf{A})\in\mathbb{R}^{n\times n}$ be continuously differentiable sufficiently many times in a region in the complex plane that contains all eigenvalues of $\mathbf{A}$. Then the function $f$ has a Fréchet derivative $L_{f}(\mathbf{A},\mathbf{E})\in\mathbb{R}^{n\times n}$ at $\mathbf{A}$ in the direction $\mathbf{E}\in\mathbb{R}^{n\times n}$ for all matrices $\mathbf{E}$ of sufficiently small norm. The Fréchet derivative satisfies
@@
THEOREM 1.
(Schweitzer [27, Theorem 2.3]) Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ and $u,v\in\mathbb{R}^{n}\backslash\{0\}$, and assume that $f$ is Fréchet differentiable at $\mathbf{A}$. Define $\mathbf{E}_{ij}=e_{i}e_{j}^{T}$, where $e_{k}=[0,\ldots,0,1,0,\ldots,0]^{T}\in\mathbb{R}^{n}$ denotes the $k$th column of the identity matrix. Then
@@
This completes the proof. ∎
@@
see, e.g., [15, p. 253]. However, when $f(t)=\exp_{0}(t)$, the computation of $f(\mathbf{A}^{T})$ requires ${\mathcal{O}}(n^{3})$ arithmetic floating point operations (flops). Therefore, the evaluation of the left-hand side of (9) demands about $8$ times more flops than the calculation of $f(\mathbf{A}^{T})$. It is cheaper to approximate $L_{f}(\mathbf{A}^{T},uv^{T})$ by using the finite-difference approximation
@@
The computation of the scalar exponential is carried out with high relative accuracy in MATLAB. However, evaluation of the matrix exponential $\exp(\mathbf{A}^{T})$ is more difficult. It can be computed in several ways; see, e.g., [15, Chapter 10] as well as [24, 30]. The accuracy achieved depends on the method use as well as on the size and properties of the matrix $\mathbf{A}^{T}$; see, e.g., [15, Chapter 10] and [24] for computed examples. We therefore include a factor $10^{3}$ in the bound $\delta$ for the relative accuracy. This bound is valid for most matrices of sizes of interest to us. Letting $\delta\approx 10^{3}\epsilon_{\text{mach}}$ with $\epsilon_{\text{mach}}\approx 2\cdot 10^{-16}$ gives $h\approx 2\cdot 10^{-4}$.   $\Box$
THEOREM 3.
@@
$\frac{f(\mathbf{A}^{T}+h\mathbf{E}_{ij})-f(\mathbf{A}^{T}-h\mathbf{E}_{ij})}{2h}$ $=$ $\mathbf{E}_{ij}+\frac{1}{2!}(\mathbf{A}^{T}\mathbf{E}_{ij}+\mathbf{E}_{ij}\mathbf{A}^{T})$
@@
The result now follows from (8) with $\mathbf{A}$ replaced by $\mathbf{A}^{T}$. ∎
The evaluation of the right-hand side of (10) with $f(t)=\exp_{0}(t)$ gives approximations of all the entries of the gradient (3). We will refer to $\|\nabla T_{\mathcal{G}}(w_{11},w_{21},\ldots,w_{nn})\|_{2}$ as the total transmission of the graph ${\mathcal{G}}$.
@@ 2.1.1 Network simplification by edge removal
Consider the cone $\mathcal{A}$ of all nonnegative matrices in $\mathbb{R}^{n\times n}$ with the same sparsity structure as $\mathbf{A}$ and let $\mathbf{M}|_{\mathcal{A}}$ denote a matrix in $\mathcal{A}$ that is closest to a given matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ with respect to the Frobenius norm. It is straightforward to verify that $\mathbf{M}|_{\mathcal{A}}$ is obtained by setting all the entries outside the sparsity structure of $\mathbf{A}$ to zero.
@@
In computations, we order the entries of ${\mathcal{E}}_{L_{f}}$ from smallest to largest and set the weights $w_{ij}$ (or $w_{ij}=w_{ji}$ if the graph is undirected) associated with the first few of the ordered edge importances to zero. Some post-processing may be necessary if the reduced graph $\widetilde{\mathcal{G}}$ obtained by removing the edges associated with the weights that are set to zero is required to be connected. A graph is said to be connected if for every pair of vertices $v_{i}$ and $v_{j}$, there is a path from vertex $v_{i}$ to vertex $v_{j}$ and from vertex $v_{j}$ to vertex $v_{i}$. Directed graphs with this property are sometimes referred to as strongly connected. A directed graph is said to be weakly connected if the undirected graph that is obtained by replacing all directed edges by undirected ones is connected. The graph $\widetilde{\mathcal{G}}$ is strongly connected if and only if the adjacency matrix associated with the graph is irreducible. This provides a computational approach to determine whether a graph is strongly connected. Further, an undirected graph is connected if the second smallest eigenvalue of the associated graph Laplacian is positive; see e.g., [11, 19].
@@ 2.1.3 Network modification by inclusion of new edges
The partial derivatives (4) reveal which edges would be important to add to a given graph to increase the communication in the network significantly, namely nonexistent edges, whose associated partial derivative is large. Let $\widehat{\mathcal{A}}$ denote the cone of the nonnegative matrices in $\mathbb{R}^{n\times n}$ with sparsity structure given by the zero entries of $\mathbf{A}$ except for the diagonal entries. The virtual importance of the nonexistent edge $e(v_{i}\to v_{j})\notin{\mathcal{E}}$ is given by the corresponding entry of $L_{f}(\mathbf{A}^{T},ee^{T})$ normalized by the total transmission. The construction of the virtual edge importance vector ^$\widehat{{\mathcal{E}}}_{L_{f}}\in\mathbb{R}^{n^{2}-n-m}$, which makes use of matrix-entries in the sparsity structure associated with $\widehat{\mathcal{A}}$, can be summarized as follows:
@@
If the graph is undirected, then the virtual importance of the virtual edge $e(v_{i}\leftrightarrow v_{j})\notin{\mathcal{E}}$ is defined as twice the corresponding entry in $L_{f}(\mathbf{A}^{T},ee^{T})$ normalized by the total transmission. Let $\widehat{\mathcal{L}}$ be the cone of the nonnegative matrices in $\widehat{\mathcal{A}}$, where all the entries in the strictly upper triangular portion are set to zero. The procedure for the construction of the virtual edge importance vector ^$\widehat{{\mathcal{E}}}_{L_{f}}\in\mathbb{R}^{(n^{2}-n-2m)/2}$, which makes use of matrix-entries in the sparsity structure associated with $\widehat{\mathcal{L}}$, becomes:
@@
In case $\|L_{f}(\mathbf{A}^{T},ee^{T})|_{\mathcal{A}}\|_{F}\ll||\nabla T_{\mathcal{G}}||_{2}$ many of large derivatives are not in the structure of $\mathbf{A}$. Then making suitable zero weights $w_{ij}$ positive, or if the graph is undirected giving suitable pairs of zero weights $\{w_{ij},w_{ji}\}$ the same positive value, might be beneficial. We recall that giving a zero weight $w_{ij}$ a positive weight $w_{ij}$ is equivalent to including a weighted edge $e(v_{i}\leftrightarrow v_{j})$ into the graph.

### v2

# SAPIENZA Università di Roma P.le Aldo Moro 5, 00185 Roma, Italy

@@ Abstract
Keywords:
Network analysis, Sensitivity analysis, Edge importance
MSC:
05C82, 15A16, 65F15
@@ 1 Introduction
Networks are helpful for modeling complex interactions between entities. A network can be represented by a graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E},\mathcal{W}\rangle$, which consists of a set of vertices or nodes $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, a set of edges $\mathcal{E}$, with $|\mathcal{E}|=m$, that connect the vertices, and a set of nonnegative edge weights $\mathcal{W}=\{w_{11},w_{21},\ldots,w_{nn}\}$. The weight $w_{ij}$ is positive if there is an edge pointing from vertex $v_{i}$ to vertex $v_{j}$; while $w_{ij}=0$ signifies that there is no such edge. Edges may be directed (and then model one-way streets) or undirected (and then model two-way streets). If the graph models a road network in which the vertices model intersections and the edges model roads, then the weights may, e.g., be proportional to the amount of traffic along each road. A network in which all positive weights are one is said to be unweighted. Descriptions and many applications of networks are provided by, e.g., Estrada Es and Newman Ne .
@@
A graph is said to be connected if for every pair of vertices $v_{i}$ and $v_{j}$, there is a path from vertex $v_{i}$ to vertex $v_{j}$ and from vertex $v_{j}$ to vertex $v_{i}$. Directed graphs with this property are sometimes referred to as strongly connected. A directed graph is said to be weakly connected if the undirected graph that is obtained by replacing all directed edges by undirected ones is connected. A graph is strongly connected if and only if the adjacency matrix associated with the graph is irreducible. This provides a computational approach to determine whether a graph is strongly connected. Further, an undirected graph is connected if and only if the second smallest eigenvalue of the associated graph Laplacian is positive; see e.g., Es ; Ne .
@@
Let $\mathbf{A}^{k}=[a_{ij}^{(k)}]_{i,j=1}^{n}$ for $k=1,2,\ldots~{}$, where $a_{ij}^{(1)}=w_{ij}$ for $1\leq i,j\leq n$. A nonvanishing entry $a_{ij}^{(k)}$ for some $k>0$ indicates that there is at least one walk of $k$ edges from vertex $v_{i}$ to vertex $v_{j}$. In case of an unweighted network, $a_{ij}^{(k)}$ represents the number of the walks of length $k$ from vertex $v_{i}$ to vertex $v_{j}$. The denominators in the terms of the expansion (1) ensure that the expansion converges and that terms $\frac{\mathbf{A}^{k}}{k!}$ with $k$ large contribute only little to $\exp_{0}(\mathbf{A})$. It follows that short walks typically are more important than long ones, which is in agreement with the intuition that messages propagate better along short walks than along long ones. This led Estrada and Rodriguez-Velazquez ER to use $\exp(\mathbf{A})$ to study properties of a graph; we use the function $\exp_{0}(\mathbf{A})$ because the term with the identity matrix $\mathbf{I}$ in the expansion of $\exp(\mathbf{A})$ has no natural interpretation in the context of network modeling. Other functions also can be used such as a resolvent or a Mittag-Leffler function; see Estrada and Higham EH and Arrigo and Durastante AD for discussions.
Estrada and Rodriguez-Velazquez ER define for graphs with a symmetric adjacency matrix the communicability matrix $\exp(\mathbf{A})$; we will use the matrix
@@
The entry $c_{ij}$, $i\neq j$, is referred to as the communicability between the vertices $v_{i}$ and $v_{j}$. A relatively large value implies that it is easy for the vertices $v_{i}$ and $v_{j}$ to communicate. Estrada and Rodriguez-Velazquez ER measure the importance of the vertex $v_{i}$ of an undirected graph by the subgraph centrality $c_{ii}+1$; we will use $c_{ii}$. Related measures of communicability can be defined when the adjacency matrix $\mathbf{A}$ is nonsymmetric; see DLCMR .
@@
where $e=[1,1,\ldots,1]^{T}\in\mathbb{R}^{n}$ denotes the vector with all entries $1$ and the superscript T stands for transposition. Benzi and Klymko BK introduced the related measure $e^{T}\exp(\mathbf{A})e$, which differs from (2) by the additive constant $n$. Note that the expression (2) is invariant under transposition. We have
@@
We are interested in investigating the importance of the edges of a graph ${\mathcal{G}}$ and, in particular, in determining which edge weights can be reduced or set to zero without significantly affecting the total communicability. A possible approach to investigate the importance of edges is to consider the line graph associated with the graph. The edges of ${\mathcal{G}}$ correspond to vertices in the associated line graph, and the importance of the edges in ${\mathcal{G}}$ can be measured by the subgraph centrality of the vertices of the line graph. This approach is investigated in DLCMR2 ; DLCMR3 . However, it is quite cumbersome to construct the line graph except for small graphs. A simple heuristic technique was proposed by Arrigo and Benzi AB , who define the edge total communicability and seek to remove edges whose removal does not reduce the edge total communicability much. This approach is quite easy to implement for networks that are small enough to allow the computation of the singular value decomposition of the associated adjacency matrix; however, it is not straightforward to use for large-scale networks. Moreover, the importance of an edge is assumed to depend on the importance of the vertices that it connects. This holds for some networks, but not for others. Therefore, this approach to edge removal may result in removals that are not in agreement with intuition. More recently an approach that uses the right and left Perron vectors of the adjacency matrix in combination with the Wilkinson perturbation to determine which weights to increase in order to increase the total communicability has been described in DLCJNR ; NR . An analogous technique is applied in EHNR to discern which weights in a weighted multilayer network can be decreased without affecting the total communicability significantly.
Another approach to study the importance of edges is to evaluate the Fréchet derivatives of the total communicability (2) with respect to the weights. This approach was advocated by De la Cruz Cabrera et al. DLCJNR and recently Schweitzer Sc described how to speed up the computations. Introduce the gradient
@@
This paper is organized as follows. Section 2 discusses how the gradient can be applied to assess the importance of edges. The first part of the section is concerned with small to medium-sized problems for which it is feasible to evaluate the gradient (3). The latter part of the section discusses the application of Krylov subspace methods to project large-scale problems to problems of fairly small size. The computations use a result by Schweitzer Sc on the evaluation of Fréchet derivatives, but differ in various aspects that speed up the computations. Section 3 reviews methods described in DLCJNR ; EHNR ; NR based on evaluating the right and left Perron vectors and the Wilkinson perturbation to determine important and unimportant edges. One of the aims of this paper is to compare the methods discussed in Sections 2 and 3. This is done in Section 4, where we also report timings. Section 5 contains concluding remarks.
We conclude this section with comments on some related methods. A scheme that combines regression, soft-thresholding, and projection is applied in DLCJR to approximate an unweighted network by a simpler unweighted network. This scheme performs well but may be expensive and is restricted to unweighted networks. Massei and Tudisco MT consider the problem of determining a low-rank perturbation $\mathbf{E}\in\mathbb{R}^{n\times n}$ to the adjacency matrix $\mathbf{A}$ so that the perturbed matrix $\mathbf{A}+\mathbf{E}$ maximizes or minimizes the robustness of the network. For instance, $\mathbf{E}$ may be chosen to maximize or minimize the trace of $f(\mathbf{A}+\mathbf{E})-f(\mathbf{A})$ for a user-specified matrix-valued function $f$. Thus, this method seeks to modify a few edge weights so that the trace is increased or decreased as much as possible. The perturbation $\mathbf{E}$ is determined by a greedy algorithm for solving an optimization problem that can be quite expensive to solve. A careful comparison with this method is outside the scope of the present paper.
@@ 2.1 Methods for small to medium-sized networks
Let the matrix function $f:\mathbf{A}\in\mathbb{R}^{n\times n}\to f(\mathbf{A})\in\mathbb{R}^{n\times n}$ be continuously differentiable sufficiently many times in a region in the complex plane that contains all eigenvalues of $\mathbf{A}$. Then the function $f$ has a Fréchet derivative $L_{f}(\mathbf{A},\mathbf{E})\in\mathbb{R}^{n\times n}$ at $\mathbf{A}$ in the direction $\mathbf{E}\in\mathbb{R}^{n\times n}\backslash\{\mathbf{0}\}$. The Fréchet derivative satisfies
@@
Theorem 2.1
(Schweitzer (Sc, , Theorem 2.3)) Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ and $u,v\in\mathbb{R}^{n}\backslash\{0\}$, and assume that $f$ is Fréchet differentiable at $\mathbf{A}$. Define $\mathbf{E}_{ij}=e_{i}e_{j}^{T}$, where $e_{k}=[0,\ldots,0,1,0,\ldots,0]^{T}\in\mathbb{R}^{n}$ denotes the $k$th column of the identity matrix. Then
@@
This completes the proof.
@@
see, e.g., (Hi, , p. 253). However, when $f(t)=\exp_{0}(t)$, the computation of $f(\mathbf{A}^{T})$ requires ${\mathcal{O}}(n^{3})$ arithmetic floating point operations (flops). Therefore, the evaluation of the left-hand side of (9) demands about $8$ times more flops than the calculation of $f(\mathbf{A}^{T})$. It is cheaper to approximate $L_{f}(\mathbf{A}^{T},uv^{T})$ by using the finite-difference approximation
@@
The computation of the scalar exponential is carried out with high relative accuracy in MATLAB. However, evaluation of the matrix exponential $\exp(\mathbf{A}^{T})$ is more difficult. It can be computed in several ways; see, e.g., (Hi, , Chapter 10) as well as RSID ; XX . The accuracy achieved depends on the method used as well as on the size and properties of the matrix $\mathbf{A}^{T}$; see, e.g., (Hi, , Chapter 10) and RSID for computed examples. We therefore include a factor $10^{3}$ in the bound $\delta$ for the relative accuracy. This bound is valid for most matrices of sizes of interest to us. Letting $\delta\approx 10^{3}\epsilon_{\text{mach}}$ with $\epsilon_{\text{mach}}\approx 2\cdot 10^{-16}$ gives $h\approx 2\cdot 10^{-4}$.   $\Box$
Proposition 1
@@
$\frac{f(\mathbf{A}^{T}+h\mathbf{E}_{ij})-f(\mathbf{A}^{T}-h\mathbf{E}_{ij})}{2h}$ $=$ $\mathbf{E}_{ij}+\frac{1}{2!}(\mathbf{A}^{T}\mathbf{E}_{ij}+\mathbf{E}_{ij}\mathbf{A}^{T})$ (11)
@@
The result now follows from (8) with $\mathbf{A}$ replaced by $\mathbf{A}^{T}$.
The above proposition is a corollary of the well-known series representation of the Fréchet derivative; see, e.g., Hi . It can be stated for an arbitrary direction matrix $\mathbf{E}$ by just replacing the last term $O(h^{2})$ in (11) with $O(\|\mathbf{E}\|_{2}\,h^{2})$. The evaluation of the right-hand side of (10) with $f(t)=\exp_{0}(t)$ gives approximations of all the entries of the gradient (3). We will refer to $\|\nabla T_{\mathcal{G}}(w_{11},w_{21},\ldots,w_{nn})\|_{2}$ as the total transmission of the graph ${\mathcal{G}}$.
@@ 2.1.1 Network simplification by edge removal
Consider the cone $\mathcal{A}$ of all nonnegative matrices in $\mathbb{R}^{n\times n}$ with the same sparsity structure as $\mathbf{A}$ and let $\mathbf{M}|_{\mathcal{A}}$ denote a matrix in $\mathcal{A}$ that is closest to a given nonnegative matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ with respect to the Frobenius norm. It is straightforward to verify that $\mathbf{M}|_{\mathcal{A}}$ is obtained by setting all the entries outside the sparsity structure of $\mathbf{A}$ to zero.
@@
In computations, we order the entries of ${\mathcal{E}}_{L_{f}}$ from smallest to largest and set the weights $w_{ij}$ (or $w_{ij}=w_{ji}$ if the graph is undirected) associated with the first few of the ordered edge importances to zero. Some post-processing may be necessary if the reduced graph $\widetilde{\mathcal{G}}$ obtained by removing the edges associated with the weights that are set to zero is required to be connected.
@@ 2.1.3 Network modification by inclusion of new edges
The partial derivatives (4) reveal which edges would be important to add to a given graph to increase the communicability of the network significantly, namely nonexistent edges, whose associated partial derivative is large. Let $\widehat{\mathcal{A}}$ denote the cone of the nonnegative matrices in $\mathbb{R}^{n\times n}$ with sparsity structure given by the zero entries of $\mathbf{A}$ except for the diagonal entries. The virtual importance of the nonexistent edge $e(v_{i}\to v_{j})\notin{\mathcal{E}}$ is given by the corresponding entry of $L_{f}(\mathbf{A}^{T},ee^{T})$ normalized by the total transmission. The construction of the virtual edge importance vector ^$\widehat{{\mathcal{E}}}_{L_{f}}\in\mathbb{R}^{n^{2}-n-m}$, which makes use of matrix entries in the sparsity structure associated with $\widehat{\mathcal{A}}$, can be summarized as follows:
@@
If the graph is undirected, then the virtual importance of the virtual edge $e(v_{i}\leftrightarrow v_{j})\notin{\mathcal{E}}$ is defined as twice the corresponding entry in $L_{f}(\mathbf{A}^{T},ee^{T})$ normalized by the total transmission. Let $\widehat{\mathcal{L}}$ be the cone of the nonnegative matrices in $\widehat{\mathcal{A}}$, where all the entries in the strictly upper triangular portion are set to zero. The procedure for the construction of the virtual edge importance vector ^$\widehat{{\mathcal{E}}}_{L_{f}}\in\mathbb{R}^{(n^{2}-n-2m)/2}$, which makes use of matrix entries in the sparsity structure associated with $\widehat{\mathcal{L}}$, becomes:
@@
In case $\|L_{f}(\mathbf{A}^{T},ee^{T})|_{\mathcal{A}}\|_{F}\ll||\nabla T_{\mathcal{G}}||_{2}$ many large partial derivatives (4) correspond to zero extra-diagonal entries of $\mathbf{A}$. Then making suitable zero weights $w_{ij}$ positive, or if the graph is undirected giving suitable pairs of zero weights $\{w_{ij},w_{ji}\}$ the same positive value, might be beneficial. We recall that replacing a zero weight $w_{ij}$ with a positive weight $w_{ij}$ is equivalent to including a weighted edge $e(v_{i}\leftrightarrow v_{j})$ into the graph.